### v1

@@ 3.4 Summary
Despite these advances in transportation studies, future research must address several key challenges. Standardizing evaluation methods is crucial to ensure consistency and comparability across studies. Future studies should also consider the dynamic characteristics of traffic data, which vary significantly across different spatial and temporal contexts. Additionally, understanding and incorporating the causal relationships within traffic data, as well as addressing ethical and privacy concerns associated with synthetic data generation, are critical areas that require further exploration. These challenges, along with other emerging trends and opportunities in the use of DGMs for transportation research, will be discussed in more detail in the following Section 5.
@@
In this section, we present practical examples of how DGMs can be applied in transportation research. To reach a broader audience, we provide two hands-on tutorials: 1) Generating Household Travel Survey Data in Section 4.1, and 2) Generating Highway Traffic Speed Contour (Time-Space Diagram) in Section 4.2. Importantly, all data and code used in these tutorials—including preprocessing scripts, model training, and inference code, and pre-trained model parameters—are available in our GitHub repository: https://github.com/benchoi93/DGMinTransportation
This tutorial aims to equip researchers with the necessary tools to explore the application of DGMs in transportation studies, rather than identifying the top-performing models within each research domain. To this end, we do not perform numerical or analytical evaluations of model performance. Instead, the focus is on providing qualitative insights, which is a common practice in DGM research. For readers interested in rigorous performance evaluations, including numerical or analytical assessments, we suggest referring to the existing literature. Nevertheless, the code provided in our repository offers a robust foundation for initiating DGM-based research, allowing researchers to extend these models for specific transportation study requirements.
@@ 4.1.1   Data and Preprocessing
In this tutorial, we applied DGMs to the 2010 Household Travel Survey data from South Korea. This dataset offers comprehensive details on household travel patterns, including trip purposes, transportation modes, and travel times across various cities and districts in the region. It serves as a valuable resource for investigating urban mobility dynamics. While the original HTS dataset covers the entire country, this study only focuses specifically on the Seoul Metropolitan Area, as illustrated in Figure 7. The Seoul Metropolitan Area encompasses Seoul, Incheon, and Gyeonggi-do located in the northwest region of South Korea. With a population of approximately 26 million, this area represents more than half of the country’s total population.
Figure 7: Seoul Metropolitan Area
The original household travel survey includes sociodemographic data and completed travel data for the households. For simplicity, we exclude the sociodemographic data and focus solely on travel data. An example of the HTS travel data used in this study is shown in Figure 8. This travel data that are composed of three different trips can be transformed into the tabular data as the Table 2. To simplify the problem, we selected the types of origin, activity, mode choice, and destination as target features for data generation. Table 3 provides a detailed description of the chosen features. In our tutorial example, each DGM is designed to generate each row that contains the data in Table 3 by learning the joint distribution of the selected features. The data includes origin and destination types divided into five categories: home, work, school, transfer points, and other locations. Transportation modes are categorized into nine discrete integers: 0 for the start of the day, 1 for the end of the day, 2 for staying at a location, and 3 through 8 representing various modes of transport such as walking, public transit, driving (driver and passenger), cycling, and taxis. Lastly, activity types include eight categories: work, education, leisure, shopping, and escort activities. It is important to note that the goal of this tutorial is not generating the complete daily travel, but rather focusing on individual trips. This allows the model to capture the relationships between features. During model evaluation, we present qualitative assessments with marginal distributions of each feature. This qualitative approach prioritizes interpretability and understanding over direct numerical performance metrics. Nevertheless, the model is trained to capture the joint probability distribution, emphasizing its ability to generate realistic combinations of attributes.
Figure 8: Example daily travel in the HTS data
Table 2: Trip chain data generated from example daily travel in Figure 8
@@ 7 Home Gyeonggi-1 08:30 PM End Stay Home Gyeonggi-1 12:00 PM
Table 3: Details of each attribute in generated data
Features Range of Value Description on Each Value
Origin Type Discrete Integer (0$\sim$5) 0: Home, 1: Work Place, 2: School 3: Transfering 4: Other
Activity Type Discrete Integer (0$\sim$8) 0: Home, 1: Travel, 2: Work, 3: Education, 4: Academy, 5: Leisure, 6: Shopping, 7: Escort, 8: Other
Mode Type Discrete Integer (0$\sim$8) 0: Start, 1: End, 2: Stay, 3: Walk, 4: Transit, 5: Car (driver), 6: Car (passenger), 7: Bicycle, 8: Taxi
Destination Type Discrete Integer (0$\sim$5) 0: Home, 1: Work Place, 2: School 3: Transfering 4: Other
@@
where $U\left[0,1\right)^{D}$ is a $D$-dimensional (data dimension) random uniform vector. This method effectively smooths the distribution and ensures stable training and generation of discrete variables.

### v3

@@ 3.4 Summary
Despite these advances in transportation studies, future research must address several key challenges. Standardizing evaluation methods is crucial to ensure consistency and comparability across studies. Future studies should also consider the dynamic characteristics of traffic data, which vary significantly across different spatial and temporal contexts. Additionally, understanding and incorporating the causal relationships within traffic data, as well as addressing ethical and privacy concerns associated with synthetic data generation, are critical areas that require further exploration. These challenges, along with other emerging trends and opportunities in the use of DGMs for transportation research, will be discussed in more detail in Section 5.
@@
In this section, we present practical examples of how DGMs can be applied in transportation research. To reach a broader audience, we provide two hands-on tutorials: 1) Generating Household Travel Survey Data in Section 4.1, and 2) Generating Highway Traffic Speed Contour (Time-Space Diagram) in Section 4.2. To keep the paper concise, detailed explanations of the model structure and its loss functions in the code are provided in Appendices F and G. Importantly, all data and code used in these tutorials—including preprocessing scripts, model training, inference code, and pre-trained model parameters—are available in our GitHub repository: https://github.com/UMN-Choi-Lab/DGMinTransportation. The tutorial code is implemented in Python, with PyTorch serving as the primary library for the tutorials. Additional requirements are detailed in the associated GitHub repository.
This tutorial aims to equip researchers with the necessary tools to explore the application of DGMs in transportation studies. Accordingly, we structured the tutorial and developed the accompanying code using one representative discrete dataset and one representative continuous dataset. Given the scope of this paper, it is impractical to address every possible data type; however, by covering these key examples, we expect to offer a wide range of valuable insights. Moreover, our goal is not to identify the top-performing models within each research domain. To this end, although we performed numerical evaluations, the results can vary significantly depending on the numerous hyperparameters and optimizers used for each model. Consequently, we want to emphasize that readers should not judge any particular model solely based on the results presented in this paper. Instead, the focus is on providing qualitative insights, which is a common practice in DGM research. For readers interested in rigorous performance evaluations, including numerical or analytical assessments, we suggest referring to the existing literature. Nevertheless, the code provided in our repository offers a robust foundation for initiating DGM-based research, enabling researchers to extend these models for specific transportation study requirements.
The code implementations for the tutorial and experiments are conducted in a Python-based environment. Key libraries used include torch (tested both on v1.13 and v2.4) for deep learning model development, leveraging GPU acceleration to handle large-scale data and neural networks efficiently, and numpy and pandas for numerical computations and data manipulation. All development is done in a notebook environment to allow for interactive data exploration, experimentation, and real-time feedback. This setup also ensures compatibility with widely used tools, including Jupyter Notebook and Google Colab, offering flexibility for development and testing.
@@ 4.1 Generating Household Travel Survey Data
Figure 7: Visual Description of Household Travel Survey dataset. (a) Study Area: Seoul Metropolitan Area (b) Example daily travel in the HTS data
@@ 4.1.1   Data and Preprocessing
Table 9: Trip chain data generated from example daily travel in Figure 7(b)
@@ 7 Home Gyeonggi-1 08:30 PM End Stay Home Gyeonggi-1 12:00 PM
In this tutorial, we applied DGMs to the 2010 HTS data from South Korea. This dataset provides comprehensive details on household travel patterns, including trip purposes, transportation modes, and travel times across various cities and districts in the region, making it a valuable resource for investigating urban mobility dynamics. While the original HTS dataset covers the entire country, this study only focuses specifically on the Seoul Metropolitan Area, as illustrated in Figure 7(a). The Seoul Metropolitan Area encompasses Seoul, Incheon, and Gyeonggi-do located in the northwest region of South Korea. With a population of approximately 26 million, this area represents more than half of the country’s total population.
The original HTS includes both sociodemographic data and completed travel data for households. For simplicity, we exclude the sociodemographic data and focus solely on travel data. An example of the HTS travel data used in this study is shown in Figure 7(b). This travel data, which is composed of three different trips, can be transformed into the tabular data as the Table 9. To simplify the problem, we selected the types of origin, activity, mode choice, and destination as target features for data generation.
In our tutorial example, each DGM is designed to generate each row that contains the data in Table 9 by learning the joint distribution of the selected features. The data includes origin and destination types divided into five categories: home, work, school, transfer points, and other locations. Transportation modes are categorized into nine discrete integers: 0 for the start of the day, 1 for the end of the day, 2 for staying at a location, and 3 through 8 representing various modes of transport such as walking, public transit, driving (driver and passenger), cycling, and taxis. Lastly, activity types include eight categories: work, education, leisure, shopping, and escort activities. It is important to note that the goal of this tutorial is not generating the complete daily travel, but rather focusing on individual trips. This allows the model to capture the relationships between features. During model evaluation, we present qualitative assessments with marginal distributions of each feature. This qualitative approach prioritizes interpretability and understanding over direct numerical performance metrics. Nevertheless, the model is trained to capture the joint probability distribution, emphasizing its ability to generate realistic combinations of attributes.